# Cyclotomic Factors and LRS-Degeneracy

### v2

@@ 2 Notation, Terminology, Preliminaries
We shall often restrict attention to square-free, content-free polynomials in ${\mathbb{Z}}[x]$: a general polynomial in ${\mathbb{Q}}[x]$ may readily be rescaled by clearing denominators and dividing out the content to yield a content-free polynomial in ${\mathbb{Z}}[x]$. We can quickly compute the radical (i.e. product of the factors in a square-free factorization) as $f/\gcd(f,f^{\prime})$ where $f^{\prime}$ denotes the formal derivative of $f$. We shall write the reverse of a non-zero polynomial as $\mathop{\rm rev}(f)\,=\,x^{d}\,f(\frac{1}{x})$ where $d=\deg(f)$.
@@
Let $k\in{\mathbb{N}}_{\geq 2}$, and let $p=1+ks$ be prime (for some $s\in{\mathbb{N}}$). Let $g$ be a primitive root modulo $p$, then $\zeta=g^{(p-1)/s}$ is a primitive $k$-th root of unity in the finite field ${\mathbb{F}}_{p}$.
@@
Let $k\in{\mathbb{N}}_{\geq 1}$. We write $\Phi_{k}$ to denote the $\bm{k}$-th cyclotomic polynomial (or cyclotomic polynomial of index $\bm{k}$), i.e. the monic polynomial whose roots are all primitive $k$-th roots of unity. It is an elementary result that $\Phi_{k}(x)\in{\mathbb{Z}}[x]$ and $\deg(\Phi_{k})=\phi(k)$.
@@ 3.1.1 Coefficient of $x^{d-1}$ and the special case $f(x)=g(x^{r})$
    Check the coefficient of $x^{d-1}$ where $d=\deg(f)$. If it is not $-1$, 00 or $1$ then $f$ is not cyclotomic. If it is 00, we take special action (described immediately below: Q4b–Q4e).
@@
    When the coefficient of $x^{d-1}$ is 00, define $r$ by saying that the second highest term has degree $d{-}r$. If the coefficient of $x^{d-r}$ is not $-1$ or $1$ then $f$ is not cyclotomic; and if $r{\not|\,}d$ or any term in $f$ has degree not a multiple of $r$ then $f$ is not cyclotomic. Otherwise we have the largest exponent $r$ such that $f(x)=g(x^{r})$.
@@
    If $\frac{d}{r}$ is odd or $\phi(\mathop{\rm rad}(r)){\not|\,}\;\frac{d}{r}$ then $f$ is not cyclotomic — here $\mathop{\rm rad}(r)$ denotes the radical of $r$, i.e. the product of all distinct primes dividing $r$.
@@
Let $f=x^{4}+x^{2}+1$. So $d=4$, and in Q4a we see that $x^{4-1}$ has coefficient 0. Q4b does not reject $f$. In Q4c we obtain $r=4-2=2$, and find that $f(x)=g(x^{2})$ for $g=x^{2}+x+1$. Q4d does not reject $f$. In Q4e we identify $g$ as $\Phi_{k}(x)$ with index $k=3$. We compute $\mathop{\rm rad}(r)=2$, and see that $\mathop{\rm rad}(r){\not|\,}k$; thus $f$ is not cyclotomic.
@@ 3.2 Testing for a cyclotomic polynomial with square-free index
The idea here is simple: we generate a list of possible candidate indexes, and then repeatedly whittle it down by computing prefixes of the corresponding cyclotomics, and keeping only those indexes whose prefixes match the highest terms in $f$. If more than one candidate index survives, we refine the list again using longer prefixes. In Algorithm 1 (CycloIndex) we write $f_{m}$ to mean the $m$-prefix of $f$, i.e. all terms of degree greater than $d{-}m$.
@@
3:$L\leftarrow\phi^{-1}(d)$ keeping only the square-free values
@@
15:if $L=[\,]$ then return not cyclotomic  end if
@@ 3.2.1 Computing an $m$-prefix
Also if $n$ has fewer than 4 odd prime factors then overflow cannot occur: if $n$ has 1 or 2 odd prime factors then the height is 1, and if there are 3 odd prime factors then the height is at most $32p/$ 3 where $p$ is the smallest odd prime factor [9].
@@ 3.3 Testing via evaluation
    Compute $V_{1}=f(1)$; this is just the sum of the coefficients. If $V_{1}\neq 1$ and $V_{1}\neq d{+}1$ then $f$ is not cyclotomic. If $V_{1}=d{+}1$: check that $d{+}1$ is prime and that ${\rm H}(f)=1$; if so, $f$ is cyclotomic with (prime) index $d{+}1$; otherwise $f$ is not cyclotomic.
@@
    Compute $V_{-1}=f(-1)$; this is an “alternating sum” of the coefficients. If $V_{-1}\neq 1$ and $V_{-1}\neq d{+}1$ then $f$ is not cyclotomic. If $V_{-1}=d{+}1$: check that $d{+}1$ is prime and that ${\rm H}(f)=1$; if so, $f$ is cyclotomic with index $2(d{+}1)$; otherwise $f$ is not cyclotomic.
@@ 3.3.1 Obtaining a candidate index (via divisibility)
A monic polynomial is cyclotomic iff it divides $x^{k}-1$ for some exponent $k>0$. The smallest such $k$ is equal to the index of the cyclotomic polynomial. We could use this as an absolute test for cyclotomicity: check whether $f(x)$ divides $x^{k}-1$ for $k=1,2,3,\ldots$ Bounds on the inverse totient function ensure that only finitely many values of $k$ need to be checked. This approach was presented in [1]. However, actually doing this is rather inefficient.
[…]
@@ 3.3.2 Supporting arguments for evaluation at 2
[…]
@@
We observe that for $r>0$ we have $2^{r}-1\neq 0$. Since $\Phi_{k}(x)$ divides $x^{k}-1$, we have that $\Phi_{k}(2)$ is non-zero and divides $12^{k}-$ 1. Moreover, since $x^{k}-1$ divides $x^{sk}-1$ for all $s\in{\mathbb{N}}_{>0}$, we immediately have that $\Phi_{k}(2)$ divides $12^{sk}-$ 1 for all $s\in{\mathbb{N}}_{>0}$.
@@ 3.3.3 Height bounds for cyclotomic polynomials
In Q5a we check the height of $f$. The outermost coefficients can be checked via A138474 as in Section 3.2.1, but once that table is exhausted we need another way: we use a second table as described here (up to degree 11 612 1591161215911\,612\,15911 612 159). If the degree of $f$ is too large for this second table, the check is skipped.
@@ 3.4 Timings
Index Factors #$\#\phi^{-1}(d)$ Time
124 525 451124525451124\,525\,451124 525 451 $8641\times 14411$ 3 35433543\,3543 354 $0.125$
120 507 533120507533120\,507\,533120 507 533 $4481\times 26893$ 6 86168616\,8616 861 $0.140$
334 482 719334482719334\,482\,719334 482 719 $5279\times 63361$ 8 14581458\,1458 145 $0.359$
399 083 849399083849399\,083\,849399 083 849 $3329\times 119881$ 10 8191081910\,81910 819 $0.406$
@@ 4.1.1 Zsigmondy’s Theorem
Let $(s_{n})_{n\in{\mathbb{N}}_{>0}}$ be a sequence of integers: let $k\in{\mathbb{N}}_{>0}$, then a primitive prime for $s_{k}$ is a prime $p$ such that $p\mid s_{k}$ and $p{\not|\,}s_{j}$ for all $j<k$. Zsigmondy’s theorem [14] tells us that if we have the sequence $s_{n}=a^{n}-b^{n}$ for coprime integers $a>b>0$ then for every index the element $s_{k}$ has a primitive prime. The only exceptions are:
@@ 4.1.2 Evaluation at a rational
Let $\beta=p/q\,\in{\mathbb{Q}}_{>1}$ with $p$ and $q$ coprime and positive. Let the sequence $\sigma_{k}=p^{k}-q^{k}$; applying Zsigmondy’s theorem we obtain a Zsigmondy factor $z_{k}$ for each index $k\in{\mathbb{N}}_{>0}$. Let the sequence $\tau_{k}=\Phi_{k}(\beta)_{num}$. Then the $z_{k}$ are Zsigmondy factors for the $\tau$ sequence — with the noted exceptions.
@@
By Lemma 2.6(a), we have $\sigma_{k}=\prod_{\kappa\mid k}\tau_{\kappa}$ for every $k\in{\mathbb{N}}_{>0}$; thus $\tau_{k}\mid\sigma_{k}$. Since each $z_{k}\mid\sigma_{k}$ and $z_{k}{\not|\,}\sigma_{\kappa}$ for any $\kappa<k$, we have that $z_{k}\mid\tau_{k}$ and $z_{k}{\not|\,}\tau_{\kappa}$ for any $\kappa{<}k$. So the $z_{k}$ are Zsigmondy factors for the sequence $\tau_{k}$. ∎
@@ 4.1.3 Algorithm
The inner “refinement” algorithm takes 3 inputs: the polynomial $f$, the evaluation point $\beta\in{\mathbb{Q}}_{>1}$, and a list of candidate indexes. It produces a refined list of candidate indexes, which might be empty. If $f$ is known to be palindromic, we set $N\leftarrow f(\beta)_{num}$ in Step 3.
@@
1:Polynomial $f$, evaluation point $\beta=p/q\in{\mathbb{Q}}_{>1}$, and $L$ a list of candidate indexes (excluding $1$ and $2$)
@@
The outer algorithm takes just the polynomial $f$ as input; it could also perform the preprocessing steps mentioned earlier. It checks explicitly for the factors $\Phi_{1}$ and $\Phi_{2}$, so that calls to the inner algorithm are not complicated by the exceptions in Zsigmondy’s theorem. It always chooses $2$ as the first evaluation point (for reasons of computational speed); thereafter it picks random small rationals. The randomness is solely to prevent an adversary from constructing a small input which causes the algorithm to perform poorly.
@@
4:Set $L_{12}\leftarrow[\,]$. If $f(1){=}0$, append $1$ to $L_{12}$. If $f(-1){=}0$, append $2$ to $L_{12}$.
@@
6:Set $L_{new}\leftarrow\textbf{RefineCandidates}(f,2,L)$triangleright see Algorithm 2
@@
9:     Set $L_{new}\leftarrow\textbf{RefineCandidates}(f,\beta,L)$triangleright see Algorithm 2
@@
If the input polynomial is monic and of even degree, we can perform, before Step 5, an initial cyclotomicity test (from Section 3) as this is very quick. The initial refinement at $\beta=2$ was quite effective in our tests: i.e. only few false positives survived. As written, Algorithm 3 has difficulty excluding the indexes $3,4$ and $6$; so in Step 10, the first time the condition $L_{new}=L$ is satisfied, we perform one more iteration but using a $\beta$ value chosen so that each of $\Phi_{k}(\beta)_{num}$ for $k=3,4,6$ has a prime factor larger than 10 0001000010\,00010 000, e.g. suitable $\beta$ values are $\frac{117}{98}$, $\frac{133}{18}$ or $\frac{169}{6}$.
@@ 4.2 Timings
$500$ $100$ 15 7311573115\,73115 731 $0.10$
1 00010001\,0001 000 $50$ 15 6961569615\,69615 696 $0.10$
1 00010001\,0001 000 $100$ 31 4873148731\,48731 487 $0.22$
1 00010001\,0001 000 $200$ 61 4446144461\,44461 444 $0.50$
We did observe a few false positives when using specially constructed square-free, palindromic polynomials such as $f=\prod_{k=2}^{201}(kx-1)(x-k)$ which have large “fixed divisors”; the fixed divisor is defined to be $\gcd\{f(n)\mid n\in{\mathbb{Z}}\}$. In all cases the false positives were indexes of cyclotomic polynomials of low degree (e.g. $3$, $4$ and $6$).
@@ 5.1 Some properties of LRS-degenerate polynomials
  • If $f\in{\mathbb{Z}}[x]$ is an even polynomial (i.e. $f(x)=f(-x)$) not of the form $a_{d}\,x^{d}$ then $f$ is $2$-LRS-degenerate.
@@
Combining these two observations, we deduce immediately that any odd polynomial (i.e. $f(-x)=-f(x)$), which is not of the form $a_{d}\,x^{d}$, is $2$-LRS-degenerate.
@@
There is a 1–1 correspondence between primitive $k$-th roots of unity and $S_{k}=\{n\in{\mathbb{N}}\mid 1\leq n<k\text{ and }\gcd(n,k)=1\}$. Namely the set of all primitive $k$-th roots is just $\{\zeta^{n}\mid n\in S_{k}\}$ where $\zeta$ is any primitive $k$-th root.
Let $\kappa$ be a factor of $k$. Let $\lambda=k/\kappa$. Then $\zeta^{i}/\zeta^{j}$ is a primitive $\kappa$-th root of unity iff $\lambda$ divides $i-j$. In particular, if $i-j=\lambda$ then $\zeta^{i}/\zeta^{j}$ is a primitive $\kappa$-th root of unity.
@@
Case $k$ is odd. For each index $j$, pick non-zero residues $r_{j}\mod p_{j}$ such that $r_{j}{+}\lambda\not\equiv 0\pmod{p_{j}}$; note that all $p_{j}>2$, so a suitable $r_{j}$ exists. Using CRT construct $n\in\{1,2,\ldots,k-1\}$ such that $n\equiv r_{j}\pmod{p_{j}}$. Then $\gcd(n,k)=1$ and $\gcd(n{+}\lambda,k)=1$. So $\zeta^{n}$ and $\zeta^{n+\lambda}$ are primitive $k$-th roots, while $\zeta^{\lambda}=\zeta^{n+\lambda}/\zeta^{n}$ is a primitive $\kappa$-th root.

### v3

@@ 2 Notation, Terminology, Preliminaries
We shall often restrict attention to square-free, content-free polynomials in ${\mathbb{Z}}[x]$: a general polynomial in ${\mathbb{Q}}[x]$ may readily be rescaled by clearing denominators and dividing out the content to yield a content-free polynomial in ${\mathbb{Z}}[x]$. We can easily compute the radical (i.e. product of the factors in a square-free factorization) as $f/\gcd(f,f^{\prime})$ where $f^{\prime}$ denotes the formal derivative of $f$. We shall write the reverse of a non-zero polynomial as $\mathop{\rm rev}(f)\,=\,x^{d}\,f(\frac{1}{x})$ where $d=\deg(f)$.
@@
Let $k\in{\mathbb{N}}_{\geq 2}$, and let $p=1+ks$ be prime (for some $s\in{\mathbb{N}}$). Let $g$ be a primitive root modulo $p$, then $\zeta=g^{(p-1)/k}=g^{s}$ is a primitive $k$-th root of unity in the finite field ${\mathbb{F}}_{p}$.
@@
Let $k\in{\mathbb{N}}_{\geq 1}$. We write $\Phi_{k}$ to denote the $\bm{k}$-th cyclotomic polynomial (or cyclotomic polynomial of index $\bm{k}$), i.e. the unique monic polynomial of smallest degree whose roots are all primitive $k$-th roots of unity. It is an elementary result that $\Phi_{k}(x)\in{\mathbb{Z}}[x]$ and $\deg(\Phi_{k})=\phi(k)$.
@@ 3.1.1 Coefficient of $x^{d-1}$ and the special case $f(x)=g(x^{r})$
    Check the coefficient of $x^{d-1}$ where $d=\deg(f)$. If it is not 1-1, 0 or $1$ then $f$ is not cyclotomic. If it is 0, we take special action (described immediately below: Q4b–Q4e).
@@
[…]
@@
[…]
@@
[…]
@@ 3.2 Testing for a cyclotomic polynomial with square-free index
The idea here is simple: we generate a list of possible candidate indexes, and then repeatedly whittle it down by computing prefixes of the corresponding cyclotomics, and keeping only those indexes whose prefixes match the highest terms in $f$. If more than one candidate index survives, we refine the list again using longer prefixes. In Algorithm 1 (CycloIndex) we write $f_{m}$ to mean the $m$-prefix of $f$, i.e. all terms of degree greater than dmd{-}m.
@@
[…]
@@
15:if L=[]L=[\,] then return not cyclotomic end if
@@ 3.2.1 Computing an $m$-prefix
Also if $n$ has fewer than 4 odd prime factors then overflow cannot occur: if $n$ has 1 or 2 odd prime factors then the height is 1, and if there are 3 odd prime factors then the height is at most $2p/3$ where $p$ is the smallest odd prime factor [9].
@@ 3.3 Testing via evaluation
    Compute $V_{1}=f(1)$; this is just the sum of the coefficients. If $V_{1}\neq 1$ and $V_{1}\neq d{+}1$ then $f$ is not cyclotomic. If $V_{1}=d{+}1$: check that d+1d{+}1 is prime and that ${\rm H}(f)=1$; if so, $f$ is cyclotomic with (prime) index d+1d{+}1; otherwise $f$ is not cyclotomic.
@@
    Compute $V_{-1}=f(-1)$; this is an “alternating sum” of the coefficients. If $V_{-1}\neq 1$ and $V_{-1}\neq d{+}1$ then $f$ is not cyclotomic. If $V_{-1}=d{+}1$: check that d+1d{+}1 is prime and that ${\rm H}(f)=1$; if so, $f$ is cyclotomic with index 2(d+1)2(d{+}1); otherwise $f$ is not cyclotomic.
@@ 3.3.1 Obtaining a candidate index (via divisibility)
An irreducible, monic polynomial is cyclotomic iff it divides $x^{k}-1$ for some exponent $k>0$. The smallest such $k$ is equal to the index of the cyclotomic polynomial. We could use this as an absolute test for cyclotomicity: check whether $f(x)$ divides $x^{k}-1$ for $k=1,2,3,\ldots$ Bounds on the inverse totient function ensure that only finitely many values of $k$ need to be checked. This approach was presented in [1]. However, actually doing this is rather inefficient.
We could make the check faster by computing the canonical remainder of $x^{r}\mod f(x)$ for $r=2d,\,3d,\,4d,\,\ldots$. Bounds on the inverse totient function show that for degrees <36 495 360<36{\,}495{\,}360 (see sequence A355667 at OEIS [8]) we can stop at $r=6d$ since $k$ cannot exceed this value. If the polynomial is indeed $\Phi_{k}$ then once we try the first exponent $r\geq k$ we obtain remainder $x^{s}$ (with s=rks=r-k), and it is easy to identify if the remainder has this form. From this remainder we easily obtain $k$ as rsr-s. However again, actually computing these polynomial remainders would still be rather costly.
Instead we mimic the above but under the mapping $x\mapsto 2$. We evaluate $f(2)$, and check that $2^{d-1}<f(2)<2^{d+1}$. Then we compute $2^{r}\mod f(2)$ for $r=2d,\,3d,\,4d,\,\ldots,7d$; to advance to the next value in the sequence we just multiply by $2^{d}$ then reduce modulo $f(2)$. Observe that on a binary computer it is easy to detect if the remainder is of the form $2^{s}$, in which case we obtain the candidate index k=rsk=r-s. We finally verify that $\deg(f)=\phi(k)$ and if so, also that $f=\Phi_{k}$ by computing $\Phi_{k}$.
@@ 3.3.2 Supporting arguments for evaluation at 2
Let $k\in{\mathbb{N}}_{>0}$ with $k\neq 6$. Let $r\in{\mathbb{N}}_{>0}$ be such that $\Phi_{k}(2)\;|\;2^{r}{-}1$. Then k|rk\;|\;r.
@@
We observe that for $r>0$ we have $2^{r}-1\neq 0$. Since $\Phi_{k}(x)$ divides $x^{k}-1$, we have that $\Phi_{k}(2)$ is non-zero and divides $2^{k}-1$. Moreover, since $x^{k}-1$ divides $x^{sk}-1$ for all $s\in{\mathbb{N}}_{>0}$, we immediately have that $\Phi_{k}(2)$ divides $2^{sk}-1$ for all $s\in{\mathbb{N}}_{>0}$.
@@ 3.3.3 Height bounds for cyclotomic polynomials
In Q5a we check the height of $f$. The outermost coefficients can be checked via A138474 as in Section 3.2.1, but once that table is exhausted we need another way: we use a second table as described here (up to degree 11 612 15911\,612\,159). If the degree of $f$ is too large for this second table, the check is skipped.
@@ 3.4 Timings
Index Factors $\phi^{-1}(d)$ Time
124 525 451124\,525\,451 $8641\times 14411$ 3 3543\,354 $0.125$
120 507 533120\,507\,533 $4481\times 26893$ 6 8616\,861 $0.140$
334 482 719334\,482\,719 $5279\times 63361$ 8 1458\,145 $0.359$
399 083 849399\,083\,849 $3329\times 119881$ 10 81910\,819 $0.406$
@@ 4.1.1 Zsigmondy’s Theorem
[…]
@@ 4.1.2 Evaluation at a rational
Let $\beta=p/q\,\in{\mathbb{Q}}_{>1}$ with $p$ and $q$ coprime and positive. Let the sequence $\sigma_{k}=p^{k}-q^{k}$; applying Zsigmondy’s theorem we obtain a Zsigmondy factor $z_{k}$ for each index $k\in{\mathbb{N}}_{>0}$ — with the noted exceptions. Let the sequence $\tau_{k}=\Phi_{k}(\beta)_{num}$. Then the $z_{k}$ are Zsigmondy factors for the $\tau$ sequence.
@@
[…]
@@ 4.1.3 Algorithm
The inner “refinement” algorithm takes 3 inputs: the polynomial $f$, the evaluation point $\beta\in{\mathbb{Q}}_{>1}$, and a list of candidate indexes. It produces a refined list of candidate indexes, which might be empty. If $f$ is known to be palindromic, we simply set $N\leftarrow f(\beta)_{num}$ in Step 3.
@@
1:Polynomial $f$, evaluation point $\beta=p/q\in{\mathbb{Q}}_{>1}$, and $L$ a list of candidate indexes (excluding 11 and 22)
@@
The outer algorithm takes just the polynomial $f$ as input; it could also perform the preprocessing steps mentioned earlier. It checks explicitly for the factors $\Phi_{1}$ and $\Phi_{2}$, so that calls to the inner algorithm are not complicated by the exceptions in Zsigmondy’s theorem. It always chooses $2$ as the first evaluation point (for reasons of computational speed); thereafter it picks random small rationals. The randomness serves to preclude the possibility of constructing a small input which causes the algorithm to perform poorly — this is standard practice in randomized algorithms.
@@
4:Set $L_{12}\leftarrow[\,]$. If f(1)=0f(1){=}0, append 11 to $L_{12}$. If f(1)=0f(-1){=}0, append 22 to $L_{12}$.
@@
6:Set $L\leftarrow\textbf{RefineCandidates}(f,2,L)$ $\triangleright$ see Algorithm 2
@@
9:  Set $L_{new}\leftarrow\textbf{RefineCandidates}(f,\beta,L)$ $\triangleright$ see Algorithm 2
@@
If the input polynomial is monic and of even degree, we can perform, before Step 5, an initial cyclotomicity test (from Section 3) as this is very quick. The initial refinement at $\beta=2$ was quite effective in our tests: i.e. only few false positives survived. As written, Algorithm 3 has difficulty excluding the indexes $3,4$ and $6$; so in Step 10, the first time the condition $L_{new}=L$ is satisfied, we perform one more iteration but using a $\beta$ value chosen so that each of $\Phi_{k}(\beta)_{num}$ for $k=3,4,6$ has a prime factor larger than 10 00010\,000, e.g. suitable $\beta$ values are $\frac{117}{98}$, $\frac{133}{18}$ or $\frac{169}{6}$.
@@ 4.2 Timings
$500$ $100$ 15 73115\,731 $0.10$
1 0001\,000 $50$ 15 69615\,696 $0.10$
1 0001\,000 $100$ 31 48731\,487 $0.22$
1 0001\,000 $200$ 61 44461\,444 $0.50$
We did observe a few false positives when using specially constructed square-free, palindromic polynomials such as $f=\prod_{k=2}^{201}(kx-1)(x-k)$ which have large “fixed divisors”; the fixed divisor is defined to be $\gcd\{f(n)\mid n\in{\mathbb{Z}}\}$. In all cases the false positives were indexes of cyclotomic polynomials of low degree (e.g. $3$, 44 and $6$).
@@ 5.1 Some properties of LRS-degenerate polynomials
  • If $f\in{\mathbb{Z}}[x]$ is of the form $g(x^{k})$ for some $k\in{\mathbb{Z}}_{\geq 2}$ but not of the form $a_{d}\,x^{d}$ then $f$ is $\kappa$-LRS-degenerate for every $\kappa\geq 2$ dividing $k$ — we note that $f$ may also be LRS-degenerate for other orders too.
@@
  • Let $f\in{\mathbb{Z}}[x]$ be divisible by $x^{d}$ for some $d\in{\mathbb{N}}$ then $f$ is $k$-LRS-degenerate iff $f/x^{d}$ is.
We observe that, for a given $f\in{\mathbb{Z}}[x]$ not of the form $a_{d}\,x^{d}$, it is computationally quick and simple to determine the largest $d,k\in{\mathbb{N}}$ such that $f(x)=x^{d}\,g(x^{k})$. If $d$ and $k$ are both even then $f$ is an even polynomial (i.e. f(x)=f(x)f(x)=f(-x)); and if $d$ is odd and $k$ even then $f$ is an odd polynomial (i.e. f(x)=f(x)f(x)=-f(-x)). We see that in both cases they are $2$-LRS-degenerate; of course, this is also immediate from the definitions of evenness and oddness.
@@
There is a 1–1 correspondence between primitive $k$-th roots of unity and $S_{k}=\{n\in{\mathbb{N}}\mid 1\leq n<k\text{ and }\gcd(n,k)=1\}$. Namely, the set of all primitive $k$-th roots is just $\{\zeta^{n}\mid n\in S_{k}\}$ where $\zeta$ is any primitive $k$-th root.
Let $\kappa>1$ be a factor of $k$, and let $\lambda=k/\kappa$. Then $\zeta^{j+\lambda}/\zeta^{j}$ is a primitive $\kappa$-th root of unity. We show how to find $j$ such that both $j$ and $j+\lambda$ are coprime to $k$, i.e. so that both $\zeta^{j}$ and $\zeta^{j+\lambda}$ are primitive $k$-th roots of unity.
@@
Case $k$ is odd. For each index $j$, pick non-zero residues $r_{j}\mod p_{j}$ such that $r_{j}{+}\lambda\not\equiv 0\pmod{p_{j}}$; note that all $p_{j}>2$, so a suitable $r_{j}$ exists. Using CRT construct $j\in\{1,2,\ldots,k-1\}$ such that $j\equiv r_{j}\pmod{p_{j}}$. Then $\gcd(j,k)=1$ and $\gcd(j{+}\lambda,k)=1$. Thus $\zeta^{j}$ and $\zeta^{j+\lambda}$ are primitive $k$-th roots, while $\zeta^{\lambda}=\zeta^{j+\lambda}/\zeta^{j}$ is a primitive $\kappa$-th root.